# Constraining the redshift of PG 1553+113 using the photohadronic model

### v1

@@ 5 SUMMARY AND DISCUSSION
Notably, the 95 per cent CL intervals derived in this study show significant overlap with redshift estimates from previous works. In some cases, narrower intervals are achieved at high confidence levels, a result of the strong statistical support provided by the extensive dataset used in the fitting process. This underscores the robustness of the photohadronic model and the reliability of the EBL corrections applied in this analysis.
@@ Acknowledgements
R. de J. P-A and G. S-C thank SECIHTI (México) for its partial support. Partial support from CSU-Long Beach is gratefully acknowledged.
@@
No new data were generated or analyzed in support of this research.
@@
$z$ = 0.38 ± 0.10 VHE spectral index – redshift correlation Malik et al. (2022)
$z$ = 0.33 ± 0.08 VHE spectral index – redshift correlation Malik et al. (2022)
@@
$z=0.3551$ and $z=0.4339$ Analysis of OVII absorbers in the line of sight of the object Nicastro (2018)
@@
Figure 1: The photohadronic model, combined with the Saldana EBL model, is applied to fit the seventeen observed VHE gamma-ray spectra of PG 1553+113. The correspoding redshift is fixed at $z=0.518$, while the normalization constants $F_{0}$ and spectral parameters $\delta$ are provided in Table 2.

### v2

@@ 5 SUMMARY AND DISCUSSION
PG 1553+113 is a luminous, highly variable gamma-ray blazar observed across the electromagnetic spectrum, from radio to TeV energies. Its extreme energy output and relativistic jet make it a pivotal target for probing particle acceleration mechanisms and jet dynamics in AGNs. As one of the brightest TeV emitters, its gamma-ray photons undergo interactions with the EBL during propagation to Earth. These interactions provide critical constraints on EBL models, which are indispensable for advancing the understanding of galaxy evolution and the cosmic infrared background. While PG 1553+113 lacks a spectroscopically confirmed redshift $z$, estimating this parameter remains vital: $z$ directly determines the source’s intrinsic luminosity and distance, enabling precise modelling of EBL photon absorption attenuating its gamma-ray emission. Such modelling is essential for testing EBL density predictions.
@@
Table 4 presents, in chronological order, the redshift constraints for PG 1553+113 reported in previous studies, alongside those derived in this work. The 95 per cent CL intervals obtained here exhibit substantial overlap with earlier redshift estimates, demonstrating consistency with prior results. Notably, narrower intervals are achieved at high confidence levels in certain cases, attributable to the robust statistical power of the comprehensive dataset utilized in the fitting process, see Fig. 4. These outcomes underscore both the reliability of the photohadronic model and the accuracy of the EBL corrections employed in this analysis. The findings confirm that the photohadronic framework accurately models all observed spectra. Among the tested EBL models, the lowest total $\chi^{2}$ value (58.82 for 58 degrees of freedom) was achieved with the Saldana-Lopez et al. (2021) model, while the Domínguez et al. (2011) model yielded the highest value ($\chi^{2}=75.67$). The 95 per cent CL intervals for the redshift’s statistical uncertainty are $0.5370.500<z<$ 0.537 (Saldana-Lopez) and $0.5270.491<z<$ 0.527 (Domínguez ).
@@ Acknowledgements
We thank the reviewer for her/his constructive remarks which helped us to improve the manuscript substantially. R. de J. P-A and G. S-C thank SECIHTI (México) for its partial support. Partial support from CSU-Long Beach is gratefully acknowledged.
@@
No new data were generated or analysed in support of this research.
@@
$z$ = 0.38 ± 0.10 VHE spectral index (lowest-flux state) – redshift correlation Malik et al. (2022)
$z$ = 0.33 ± 0.08 VHE spectral index (average) – redshift correlation Malik et al. (2022)
@@
$0.4339\leq z$ Analysis of intervening OVII absorbers along the line of sight of the object Nicastro (2018)
$0.3551\leq z$ Analysis of intervening OVII absorbers along the line of sight of the object Nicastro (2018)
@@
Figure 1: The photohadronic model, combined with the Saldana EBL model, is applied to fit the seventeen observed VHE gamma-ray spectra of PG 1553+113. The corresponding redshift is fixed at $z=0.518$, while the normalization constants $F_{0}$ and spectral parameters $\delta$ are provided in Table 2.
@@
Figure 4: Graphical comparison of historical redshift estimates for PG 1553+113, which complements the quantitative constraints presented in Table 4. The figure provides a consolidated overview of the redshift ranges reported in prior studies alongside those derived in this work.